i have tried this problems plenty of times and still cant figure out the right answer. not a graph equation
1-2(3c-5)=10-6c
and
5(b-1)-3b=7b-5(1+b)

Found 2 solutions by checkley77, scott8148:
Answer by checkley77(12844)

1-2(3c-5)=10-6c
1-6c+10=10-6c
-6c+6c=10-1-10
0c=-1 no solution for c.
-----------------------------
5(b-1)-3b=7b-5(1+b)
5b-5-3b=7b-5-5b
5b-3b-7b+5b=5+5
0b=10 no solution for b.

Answer by scott8148(6628)

1 - 6c + 10 = 10 - 6c ___ NO SOLUTION



5b - 5 - 3b = 7b - 5 - 5b
2b - 5 = 2b - 5 ___ NO UNIQUE SOLUTION